Can you show me how to solve this problem please?
8(2x-9)-3=8(x-8)+(5)

Answer by uma(370)

8(2x-9) - 3 = 8(x-8) + 5
First we remove the parenthesis by multiplying the terms inside the parenthesis and the outside factor.
==> 16x - 72 - 3 = 8x - 64 + 5
==> 16x - 75 = 8x - 59 [simplifying the values]
==> 16x - 8x = -59 + 75 [Adding 75-8x to both the sides ]
==> 8x = 16 [simplifying the like terms]
==> 8x/8 = 16/8
==>x = 2
Good Luck!!!